Help
A series RLC circuit has components with the follo
Answer #1

Given Data
         L = 16.0mH
           = 16.0*10-3H
        C = 82 nF
           = 82*10-9 F
        R = 15.0
       and Vmax= 100 V

a) Resonant Frequency, fr = 1/(2*π*√(L*C))

   Resonant Frequency, fr = 1 / [2*π*√(16.0*10-3*82*10-9 )]

                                       ≈ 4394 Hz = 4.394 kHz

b) I max = Vmax / R

I max = 100 / 15

              = 6.67 A

c) Q = L*ω0 / R

Here ω0 = 1/ √(LC)

ω0 = 27608

Q = L*ω0 / R

Q = (16.0*10-3)*27608 / 15

         ≈ 29.45 Quality factor

d) VL = XL *I max

           = ω0*L*Imax

           = (27608 )*(16.0*10-3)*6.67

           ≈ 2946 V = 2.946 kV